OpenStudy (anonymous):

HELP! how do you prove sin^-1 = cscx ?

OpenStudy (lgbasallote):

\[\sin ^{-1} x \ne \csc x\] however if you meant \[(\sin x)^{-1} = \csc x\] then that's right

OpenStudy (lgbasallote):

\[(\sin x)^{-1} \implies \frac{1}{\sin x} \implies \csc x\] does that help?

OpenStudy (jiteshmeghwal9):

\[\frac{1}{\sin}=cscx\]

OpenStudy (lgbasallote):

@best.shakir what just happened? where did 1-cos x come from?

OpenStudy (anonymous):

WAIT if csc^-1 = sin^-1 (1/x) AM i correct?

OpenStudy (anonymous):

sorry

OpenStudy (lgbasallote):

...that doesnt sound right

OpenStudy (lgbasallote):

@lsugano

OpenStudy (jiteshmeghwal9):

having the power of -1 the number will come as reciprocal &reminding the property that \[1/\sin=\csc\]

OpenStudy (lgbasallote):

\[(\csc x)^{-1} \implies \sin x \ne \sin (\frac 1x)\]

OpenStudy (anonymous):

@lgbasallote are they equal: \[\large \sin^2(x) = (\sin(x))^2\]

OpenStudy (anonymous):

no im talking about identities for the inverse functions! so its not an exponent!

OpenStudy (lgbasallote):

@waterineyes yes

OpenStudy (lgbasallote):

i know @lsugano

OpenStudy (lgbasallote):

do you mean arc cosecant?

OpenStudy (anonymous):

Are they equal ?? \[\large \sin^{-1}x = (\sin(x))^{-1}\]

OpenStudy (lgbasallote):

no

OpenStudy (anonymous):

How ??

OpenStudy (unklerhaukus):

\[\sin(x)=\frac{ \text{opposite}}{\text{hypotenuse}}\] \[\csc(x)=\frac{\text{hypotenuse}}{ \text{opposite}}\]

OpenStudy (lgbasallote):

\[\sin^{-1} x\] that is arcsin \[(\sin x)^{-1} \implies \frac{1}{\sin x} \implies \csc x\]

OpenStudy (jiteshmeghwal9):

u can prove it as since sinx=p/h\[\color{red}{{1\over \frac{p}{h}}={h \over p}}\]

OpenStudy (jiteshmeghwal9):

h/p=csc

OpenStudy (lgbasallote):

anyway..i have to go now...good luck on all of you because you'll need it considering how vague the question is ;)

OpenStudy (anonymous):

You have just replaced 2 by -1 and you are saying no ??

OpenStudy (anonymous):

nonono i havent reached arc cosecant yet. the book said that if x≤-1 or x≥1 then csc^-1x= sin^-1 (1/x) i want to know the opposte!!

OpenStudy (anonymous):

im so lost....

OpenStudy (jiteshmeghwal9):

first of all u have to know that sin=p/h

OpenStudy (jiteshmeghwal9):

sin^-1=1/sin

OpenStudy (jiteshmeghwal9):

is it clear to u @Isugano ??????????/

OpenStudy (anonymous):

yes okay

OpenStudy (jiteshmeghwal9):

1/p/h=